OpenStudy (anonymous):

Both solutions are extraneous because they don't satisfy the original equation. What error did Eloise make?

OpenStudy (anonymous):

here is the question:

OpenStudy (phi):

any thoughts ?

OpenStudy (anonymous):

i tried solving on my own but i got stuck and this part: \[(\sqrt{-2x})^2 = (x+2)^2\]

OpenStudy (anonymous):

how would i square x? do i square both altogether?

OpenStudy (anonymous):

im talking about the (x+2)^2 part

OpenStudy (phi):

There is a mistake before that, so let's not look at that. Instead, look at \[ \sqrt{-2x+1}-1 \] remember, you can think of square roots as exponents, like this: \[ (-2x+1)^\frac{1}{2} -1 \] you can't subtract off the 1 inside the parens.

OpenStudy (phi):

if you have a calculator you will find \[ \sqrt{5}-1 \ne \sqrt{4} \]

OpenStudy (phi):

the last good step is \[ \sqrt{-2x+1}= (x+3) \] and what we do next is square both sides what do we get on the left side after we square it ?

OpenStudy (phi):

squaring a square root "gets rid" of the square root sign

OpenStudy (anonymous):

why can't i subtract the one? and what did you do to put in the calculator to get that and how does that tell you it's incorrect to subtract the one? how would i know if what i subtract is the right thing to do or not? (a lot of questions sorry, i just want to learn the subject well) on the left side i got: -2x +1 i couldn't solve the right side

OpenStudy (phi):

Did you learn order of operations ?

OpenStudy (anonymous):

yes! PEMDAS right?

OpenStudy (phi):

ok, so you know (for example) 2*2 - 1 you do 2*2 first to get 4 and then 4-1 does that make sense ?

OpenStudy (phi):

and you do *not* do the 2-1 first: 2*2-1 becomes 2* 1 = 2 <-- wrong.

OpenStudy (anonymous):

yes that makes sense!

OpenStudy (phi):

The point is, you are only allowed to add or subtract after you first do parens, exponents and multiply or divide. (pretty low on the totem pole) square roots (and *everything* inside them) are done first so for example \[ \sqrt{4} -1 \] is not \[ \sqrt{4-1} \] those are two different things

OpenStudy (phi):

I hope you don't have to solve Eloise's problem because it requires using the quadratic formula, and if you don't know how to do (x+3)^2 you probably don't know the quadratic formula ... though perhaps it's in your notes?

OpenStudy (anonymous):

oh i think i understand, to subtract from the one you would just have to square it first to get: -2x +1 i don't have to solve her problem but i want to anyway, you don't have to help me with the rest of it if you don't want. but thank you thus far!

OpenStudy (phi):

so far you have \[ \sqrt{-2x+1}= (x+3) \] square both sides: \[ -2x+1= (x+3)^2 \] remember, (x+3)^2 means (x+3)(x+3) if we multiply that out (see videos) we get \[ x^2+6x+9\] so the problem is \[ -2x +1 = x^2+6x+9\] next add +2x to both sides \[ 2x -2x + 1 = x^2 +6x+2x+9\] on the left side 2x's take away 2x's is 0 on the right side, 6 x's plus 2 more x's is 8 x's we get \[ 1 = x^2+8x+9\]

OpenStudy (phi):

next, add -1 to both sides \[ 1+-1 = x^2+8x+9+-1\] on the left side we get 1-1=0 on the right, simplify 9-1 to get 8 we have \[ 0 = x^2+8x+8\] which is usually written this way: \[ x^2+8x+8=0\]

OpenStudy (phi):

that equation is solved using the quadratic formula https://www.khanacademy.org/math/algebra/quadratics/solving-quadratics-using-the-quadratic-formula/v/using-the-quadratic-formula which takes a bit of time to understand and use.

OpenStudy (phi):

the two answers you get are \[ -4 + 2\sqrt{2} \] and \[ -4 -2 \sqrt{2} \] the 2nd solution is *extraneous* and we need a calculator to evaluate these messy numbers.

OpenStudy (anonymous):

oh okay, i think i remember a bit of the formula! could you not simplify those expressions into \[-2\sqrt{2}\] and \[-6\sqrt{2}\]?

OpenStudy (phi):

no, the \[ 2 \sqrt{2} \] means 2 times sqr(2) i.e. multiplying two numbers in other words that is like doing the -4+2 first in -4 + 2*2 to get -2*2 = -4 but we are not allowed to do that

OpenStudy (phi):

remember add/subtract is very low priority, and multiply comes first

OpenStudy (anonymous):

oh right! due to pemdas, right!

OpenStudy (phi):

if you see two numbers being multiplied, you can't subtract or add anything to either of them. 1+2*3 if you added the 1 to the 2 first, we would get 3*3 but we know we have to do 2*3 first and the answer is 1+6=7

OpenStudy (phi):

so try to think like this: square undoes square root divide undoes multiply subtract undoes add (and vice versa)

OpenStudy (anonymous):

oh this is making so much sense to me thank you so much! i'm gonna try to solve the rest of the problem using the two answers, but thank you for helping me understand the ins and outs because I overlook them so often and make "silly mistakes"! thank you again